A man throws a stone in upward vertical direction from ground and reach Max height H. During ascent stone travels a distance H/5 in the last second. Then H is equal to :?

Grade:11

1 Answers

Saurabh Koranglekar
3 years ago
Dear student

s(t) – s(t-1) = u(t-(t-1)) + 0.5g ( t^2 – (t-1)^2 )

find t and substitute

in v^2 = u^2 + 2gS

Regards